### v4

@@ 4.1 Baseline Methods
Table 4: Regression task results. Mean Squared Error (MSE) and Mean Absolute Error (MAE).
@@
4.2 Classification Task
4.2.1 Implementation Details
We conducted two classification experiments: binary classification, in which we classify only struggle and non-struggle, and four-class classification, in which we predict the level of struggle. In either case, the model predicts the probability for each struggle level and is supervised by the Cross-Entropy (CE) loss. We use top-1 accuracy as an evaluation metric.
During training, we uniformly sample 128 frames from each video. The input frames are randomly cropped from $256\times 456$ to $256\times 256$. We also implement a horizontal flip at a probability of 0.5 during the training stage. In the test stage, we perform 5 random crops and ensemble the predictions. For TSN [58], we implement the BN-Inception base architecture with the average consensus for the output from each of the segments. We use ImageNet [45] pre-trained features and only train the fully-connected layer for classification. To overcome overfitting, we set the dropout rate to 0.5. We train for 65 epochs using mini-batch stochastic gradient descent with a momentum of 0.9 and weight decay 5e-4. The learning rate starts from 0.0001 and decreases 10×10$\times$10 × in epochs 30 and 60. For the SlowFast model and its variants we use the SlowFast-R50 and load the parameters pre-trained on Kinetics-400 [25]. During training, we freeze the parameters in the backbone and train the rest of the MLP layers. We use the same optimizer but the weight decay is set to 1e-5. We train these deep models for 70 epochs and implement a cyclic cosine decay learning rate scheduler with fixed cycle restart intervals [2] in which the learning rate starts from 0.001 and decreases 100 times in 50 epochs and followed by two restart intervals for 10 epochs each with the restart learning rate increase 10 times on the minimum value and decrease back to the minimum in every 10 epochs.
@@ 4.2.1 Implementation Details
Table 5: Label distribution learning results. Mean Absolute Error (MAE) and Spearman’s Rank Correlation (Spearman’s Rho).
@@
4.2.2 Classification Results
The classification task results are shown in Table 3. The first three rows are naive baselines: (1) random performance using the frequency of classes as the probability; (2) majority class in percentage; and (3) voter’s baseline in which we use the mode of the voters’ to calculate the averaged accuracy compared with the expert annotations. In general, TSN has the lowest accuracy, because it doesn’t consider temporal relationships compared to other models based on 3D-CNNs. We can also see that there remains a gap between the voters’ baseline accuracy and the best performance of our deep models. We highlight the importance of temporal modelling in further ablations in Section 5. Finally, we note that the three 3D-ConvNets have similar accuracy, which means there is little benefit gained from implementing additional MLP/Transformer-based architectures for integrating temporal information in comparison with simply using global average pooling (GAP) over the temporal sampled segments. We hypothesise this is because we evaluate our models on short video sequences so the potential power of the MLP/Transformer layers is limited.
4.3 Regression Task
4.3.1 Implementation Details
In the regression task, the labels for struggle are treated as continuous scores indicating the degree of struggle, ranging from 1 (definitely non-struggle) to 4 (definitely struggle). We add a fully connected layer that takes the feature vectors as input and outputs a regression score. The predicted regression score is supervised by Mean Squared Error (MSE) together with target regression labels. The evaluation metrics include Mean Squared Error (MSE) and Mean Absolute Error (MAE) for regression and top-1 binary or four-way classification accuracy for classification. The hyperparameter settings of the optimization process are the same as those in the classification task.
4.3.2 Regression Results
Regression results are shown in Table 4. To obtain the classification results, we quantize the regression score into binary and four-way classification labels using evenly divided thresholds and then calculate classification accuracy. VTN-SlowFast-ViT generally has the lowest MSE and MAE and the highest classification accuracy rate on activities pipes and towers, followed by SlowFast-R50 and then TSN. We find that there is a greater variability of model performance when using continuous labels compared to the classification task. However, this trend does not hold on the Tent-Struggle dataset where the model VTN-SlowFast-ViT does not outperform the SlowFast-R50, either on regression metrics or on the converted classification metrics. This may be because the dataset tent is the most complex for struggle determination—it involves outdoor tent assembly scenes with five different sub-actions, so it is hard to optimize the extra ViT layer in the model VTN-SlowFast-ViT by only supervising a regression output.
4.4 Label Distribution Learning Task
4.4.1 Implementation Details
@@
In label distribution learning, we train a deep model to predict voters’ struggle label frequency distribution given video samples. In our implementation, the final layer output of the model has four nodes corresponding to each of the four struggle levels ranging from 1 (definitely non-struggle) to 4 (definitely struggle) followed by a SoftMax layer so that each node’s output value ranges from 0 to 1 with the summation equaling to one. We optimize the Kullback–Leibler (KL) divergence loss function to align the predicted distribution with the frequency distribution of the voters’ struggle labels for each video sample. For evaluation, we calculate Spearmans’ Rank Correlation (Spearmans’ Rho) between the predicted frequency distribution and the voters’ struggle label frequency distribution so that the value ranges from -1 to 1 with a higher score indicating that the predicted distribution better aligns with its target. We also calculate the Mean Absolute Error (MAE) between the predicted distribution and the target distribution. For converted classification, we calculate the accuracy rate between the predicted highest frequency label and the voters’ mode label. The hyper-parameter settings are the same as those in the previous tasks.
Figure 5: Importance of Multiple Frames. The three bar charts on the left depict a binary classification scenario for the Pipes-Struggle, Tent-Struggle, and Tower-Struggle activities, while the three bar charts on the right depict a four-way classification scenario in the same order. Each group of bars represents the base model (blue); evaluating using a single frame (orange); and training and evaluating using a single frame (green). Error bars show the standard deviation (StdDev) across the sampled frames from three proportions (25%, 50%, and 75%) of the video frames counted starting at the beginning.
Figure 6: Importance of Temporal Ordering. The three bar charts on the left depict a binary classification scenario for the Pipes-Struggle, Tent-Struggle, and Tower-Struggle activities, while the three bar charts on the right depict a four-way classification scenario in the same order. Each group of bars represents the input frames in the correct order (blue); the evaluating using shuffled frames (orange); and the training and evaluating using shuffled frames (green).
4.4.2 Distribution Results
The results for label distribution learning are shown in Table 5. Due to the continuous nature of MAE and Spearman’s Rho evaluation metrics, it is possible to demonstrate discrepancies in the evaluation metrics between the model TSN, SlowFast-R50, and VTN-SlowFast-ViT that lead to the same trend of discrepancies in the corresponding converted accuracy rate. On all evaluation metrics we implemented in this task, the 3D-ConvNet SlowFast outperforms the 2D-ConvNet TSN, and the VTN-SlowFast-ViT is typically superior to the SlowFast. The spatial-temporal features of SlowFast appear to be more useful than the 2D spatial features in TSN, and the additional ViT layer provides a better fit for the mode label distributions when compared to vanilla SlowFast-R50. The classification results shown in Table 5 are obtained from the predicted voters’ label frequency distribution. we convert the frequency distribution into class labels by selecting the label with top-1 frequency as a four-way classification label and combining the predicted frequency of 1 with 2, and 3 with 4 to get binary classification frequency for binary classification. Therefore, the converted classification results can also be considered a classification baseline based on voters’ mode labels.
5 Ablation Study
In this section, we investigate the importance of temporal modelling and motion understanding for the struggle determination task. Specifically, we conduct two experiments: Firstly, how much information can be gleaned from a single frame, and secondly, how important is the temporal ordering of frames.
5.1 Multiple Frames
To show the necessity of using multiple frames, we conducted experiments using one frame and made a comparison with the original multi-frame setting. We first loaded the model parameters trained using 128 frames which is the standard setting for our experiments and then ran only the test stage by loading one frame from each of the video sequences. The one frame is selected from proportions of 25%, 50%, and 75% respectively from the beginning of each video segment. Note that we duplicated the sampled one frame to meet the minimum number of input frames required by 3D-ConvNets, which is 8 for SlowFast, 128 for SlowFast-gMLP and VTN-SlowFast-ViT. To maintain consistency between the training and testing stages, we conducted additional experiments by randomly selecting one frame from each video during the training stage and one frame from the same proportions during the testing stage. The results for the two cases described above are shown in Figure 5. We report the one-frame test accuracy using average accuracy from proportions of 25%, 50%, and 75% respectively. There is a significant decrease in accuracy across all of the models we tested on all three activities. And because the video samples in the dataset are randomly trimmed into 10-second segments, the accuracy rates for the frame sampled from 25%, 50%, and 75% are expected to be similar, which means a small standard deviation (StdDev) of accuracy close to 0 in the bar plots. Even though we trained the model using a single frame, the test results still showed a notable decrease in accuracy, despite a slight increase compared to testing the models with only one frame. These results indicate that no sufficient visual information is provided for the deep learning models to rely on just one frame to determine whether a person is struggling. Therefore, effectively capturing the useful temporal motion features is crucial for our struggle determination task. A sequence of frames can typically reveal anomalous hand motions such as ‘unsmooth’ motions, stopping, and repeating that indicate the person did not have things under control. A single frame cannot represent these features. To test this idea more rigorously, however, it is important to present the models with a sequence of frames that do not contain useful temporal motion features.
@@ 5.1 Multiple Frames
5.2 Temporal Ordering
We ran a shuffle-frame test to show the importance of temporal ordering. We did not consider TSN because it only considers 2D spatial features and calculates the average of the output logits of all the input frames, making it insensitive to the order of the frames. We first loaded the model parameters pre-trained using 128 frames in the right order. Then we loaded the same number of 128 frames from each of the video samples and shuffled them randomly so that we disrupted the correct order of the video frames. We ran the model in inference mode ten times and calculated the average accuracy. As shown in Figure 6, the accuracy rate decreased significantly to the random level, which means the models are very sensitive to the frame orders. To make a fair comparison, we conducted another series of experiments in which we also trained the models using shuffled frames and then reported the test shuffle-frame accuracy. Figure 6 shows that after training using shuffled frames, the models can perform much better than when solely conducting the test stage, but still lower than the results when using the frames in the correct order. Our illustration suggests that the deep models learned key temporal dependencies of video sequences when they were trained using the correct order of frame sequences; as a result, when we test the models using shuffled frames, their accuracy approaches random performance.
@@ 6 Attention Visualization
In this section, we provide visualization towards what the models are focusing on via their attention weights/activation maps [5]. This is key to ensuring that the models are learning how to detect the struggles of participants. Specifically, we visualize the activation maps from the backbone and the attention scores in the self-attention layer of the VTN-SlowFast-ViT model. In Figure 7 we show the visualization heatmaps from the final convolutional layer in the backbone network using Grad-CAM++ [5] and the attention score visualized using green bars. We present one sample from each of the three activities: Pipes-Struggle, Tent-Struggle, and Tower-Struggle. We can observe that the model chooses to focus on the hands and active objects in the frames, despite occasionally being noisy and focusing on the background. The corresponding green bars below show that the model is giving each frame varying amounts of attention. For example, in the Pipes-Struggle visualised frames, the heatmaps highlight the hand of the person holding the pipes as well as the pipes that are lying in the background. Similar to this, in the Tent-Struggle examples, the model focuses on the tent’s poles and the hands of the person attempting to put it up, as shown by the heatmaps. In the Tower-Struggle examples, the heatmaps mainly lie on the hands and the Hanoi puzzle blocks. Due to the limit of the 10-second video clips that cannot feature the entire action, the green bars are typically not aligned with explicit actions but still tend to indicate a preference for specific frames within the sampled video sequence.
@@ 7 Discussion and Conclusions
In this paper, we have shown and discussed how struggle determination is an important component of human video understanding and a key element for building assistive systems. We have introduced a dataset including three activities for the struggle determination task, with a total of 5.09 hours of video.
Our results reveal opportunities for improvement in the four-way classification task aimed at finer-grained assessment of struggle. The similar accuracy rates between four-way to binary conversion and binary classification indicate that the main challenge lies in distinguishing the degree of struggle. This is key for useful and dependable assistive systems. The tent assembly task has the lowest accuracy ($40\%$), for the four-way task, supporting our hypothesis that the ambiguity on task judgement for the tent pitching activity is harder than for the other activities that have more structure. The regression and label distribution learning tasks have encouraging results, demonstrating their efficacy in distinguishing model differences and determining struggle in different ways. In particular, the VTN-SlowFast-ViT model overall performs well in the label distribution learning tasks, yet albeit better than random, it still achieves low performance ($34\%$) for the four-way Tent-Struggle.
Several limitations and challenges remain: (1) Limited background scene variations across the dataset, consisting mostly of two indoor scenes and one outdoor scene for tent assembly. (2) We have only video-level ‘weak’ labels for struggle, finer temporal localisation and further task knowledge could help the model to improve performance. (3) Current pre-trained model parameters based on coarse-grained video action recognition may be sub-optimal for fine-grained struggle determination. (4) By using randomly trimmed 10-second video segments, we may overlook certain types of struggling that are more obvious in the entire video and can only focus on ‘low-level’ struggling detection. (5) A better understanding and validation of what is an appropriate human (or statistical) baseline is likely needed to validate model performance.
In conclusion, this paper aims to motivate and support further work on determining the level of struggle in video. A new annotated dataset for struggling with three real daily living activities is introduced. This data is with video-level annotations from both crowd voters and human experts. Three research challenges, namely classification, regression, and label distribution learning, are presented along with definitions and baseline results using various deep architectures. We hope this research provides valuable insights for additional work on fine-grained egocentric video understanding and supports new methods and directions to benchmark assistive systems.
@@
Dataset Access Our struggle determination dataset can be found via this URL:
https://github.com/FELIXFENG2019/Struggle-Determination. (Temporary during review, archival repository will be provided at acceptance.)
@@
Acknowledgments We are grateful to Prof. Dima Damen’s contributions for early discussions and involvement in the initial investigation of this work. Shijia Feng is supported by a scholarship from the China Scholarship Council (No.202109210007). Initial research, data capture and annotation supported by EPSRC grant GLANCE (EP/N013964/1).
@@ Appendix A Related Datasets
In comparison, our struggle determination datasets encompass three tasks that demonstrate different types of skills in assembling pipes, pitching tents, and playing the Tower of Hanoi game. These datasets cover indoor and outdoor scenes and include both rigid and non-rigid objects. Additionally, we specifically use a struggle scoring scale, and thus, the annotations in our dataset are not limited to binary labels but are further fine-grained to indicate the degree of struggle. Aside from the focus on struggle, this distinguishes them from the class labels in AR, action quality scores in AQA, ranking labels in skill determination, and even the binary contrastive labels in VAD. Hence, it is essential to create a dataset with a focus on struggle annotations to train deep models for determining struggle and quantifying its degree.
@@ B.1 Instructions for Participants
During our data collection process, instructions were given to our in-person participants both verbally and written on paper. For the plumbing pipes task, there are two instruction diagrams illustrating how to assemble the pipes correctly, one version is easy to follow, shown in Figure 8, and another more difficult version, shown in Figure 9 so that the participants using a hard version of instructions are more likely to show signs of struggle. For the tent-pitching task, we used the paper instructions from the manufacturer as shown in Figure 10. For the Tower of Hanoi game task, the researchers only provided the instructions verbally introducing the basic rules of the game, such as only one disk can be moved at a time and a larger disk cannot be placed on top of a smaller one.

### v5

@@ 5.2 Temporal Ordering
Table 6: Struggle regression results. Mean Squared Error (MSE) and Mean Absolute Error (MAE), along with converted binary and four-way classification accuracy.
@@
MViTv2 [31] 0.955 0.828 81.29% 33.41%
To further examine the role of temporal motion features in determining struggle, we conducted a shuffle-frame test to assess the importance of frame ordering. In this experiment, we excluded TSN [68], as it only considers 2D spatial features and averages the output logits of all input frames, making it inherently insensitive to frame order.
We began by loading model parameters pre-trained on 128 frames in their correct sequential order. Then, for each video sample, we randomly shuffled the same number of 128 frames, disrupting their natural temporal sequence. The model was run in inference mode ten times, and we reported the average accuracy across these runs.
As shown in Figure 7, the accuracy dropped significantly to a near-random level, indicating that the models for determining struggle are highly dependent on the correct frame order. To ensure a fair comparison, we conducted an additional experiment in which we trained the models using shuffled frames and then evaluated them under the same conditions. While training on shuffled frames led to improved performance compared to testing on shuffled frames alone, the accuracy remained lower than when the models were trained and tested on correctly ordered frames.
These results suggest that deep learning models learn crucial temporal dependencies when trained on properly ordered sequences to determine struggle. When this temporal structure is disrupted, model performance deteriorates, confirming that the models rely on motion continuity to infer struggle-related cues.
5.3 Alternate Training Strategies
This section explores alternative training strategies for struggle determination, focusing on struggle regression and struggle label distribution learning. Struggle regression provides a full regression baseline, complementing the regression-to-classification approach. Struggle label distribution learning trains models to predict the distribution of voters’ annotations, making better use of crowd-sourced struggle labels. We detail the implementation of these strategies and evaluate their effectiveness across different deep models.
5.3.1 Struggle Regression
Implementation Details
In the regression task, struggle labels are treated as continuous scores ranging from 1 (definitely non-struggle) to 4 (definitely struggle), representing the degree of struggle. A fully connected layer is added for regression, which takes the feature vectors as input and outputs a regression score. The predicted score is supervised using Mean Squared Error (MSE) with the target regression labels. Evaluation metrics include MSE and Mean Absolute Error (MAE) for regression, as well as top-1 accuracy for binary and four-way classification. To ensure a solid conclusion on whether a person is struggling and to what extent, the continuous regression scores are quantized into struggle-level categories by thresholds when computing classification accuracy for the struggle regression model. The hyperparameter settings for optimization remain the same as in the struggle classification task.
Regression Results
Regression results are shown in Table 6. Overall, VTN-SlowFast-ViT achieves the lowest MSE and MAE and the highest binary and four-way classification accuracy for the Pipes-Struggle and Tower-Struggle activities. SlowFast-R50 [16] follows closely, showing competitive regression performance and classification accuracy, particularly for Tent-Struggle. This suggests that incorporating a vision transformer layer on top of the frame-level features from the SlowFast [16] backbone as a temporal model improves the fit to the struggle regression labels. In contrast, TSN [68] and MViTv2 [31] generally underperform compared to these two models.
5.3.2 Struggle Label Distribution Learning
@@
Table 7: Struggle Label distribution learning results. Mean Absolute Error (MAE) and Spearman’s Rank Correlation (Spearman’s Rho), along with the converted binary classification accuracy and four-way classification accuracy.
@@
MViTv2 [31] 0.14 0.5264 80.27% 50.81%
Implementation Details
@@
In label distribution learning, we train a deep model to predict the frequency distribution of voters’ struggle labels given video samples. In our implementation, the final layer output of the model has four nodes corresponding to each of the four struggle levels ranging from 1 (definitely non-struggle) to 4 (definitely struggle), followed by a SoftMax layer so that each node’s output value ranges from 0 to 1 with the summation equalling to one.
We optimize the Kullback–Leibler (KL) divergence loss function to align the predicted distribution with the frequency distribution of the voters’ struggle labels for each video sample. For evaluation, we calculate Spearmans’ Rank Correlation (Spearmans’ Rho) between the predicted frequency distribution and the voters’ struggle label frequency distribution so that the value ranges from -1 to 1, with a higher score indicating that the predicted distribution better aligns with its target. We also calculate the Mean Absolute Error (MAE) between the predicted and the target distribution.
The converted classification accuracy is derived from the predicted voters’ label frequency distribution. To convert the frequency distribution into class labels, we select the top-1 frequency label for four-way classification. For binary classification, we merge the predicted frequencies of labels 1 and 2 into the “non-struggle” category and labels 3 and 4 into the “struggle” category. We determine the accuracy by comparing the predicted highest-frequency label with the voters’ mode label.
The hyperparameter settings remain consistent with those used in previous tasks.
Distribution Results
The results for label distribution learning are shown in Table 7. Due to the continuous nature of MAE and Spearman’s Rho evaluation metrics, discrepancies can be observed among the implemented baseline models. These discrepancies follow the same trend when converted to accuracy rates.
The MViTv2 [31] model achieves the highest binary and four-way classification accuracy rates on Tent-Struggle and Tower-Struggle. However, it does not perform as well in struggle label distribution regression, as indicated by its lower Spearman’s Rho and MAE scores. In contrast, VTN-SlowFast-ViT consistently achieves the best label distribution regression performance, with the lowest MAE and highest Spearman’s Rho across all activities. The converted classification accuracy is also among the highest although slightly lower than that of MViTv2 [31] on Tent-Struggle and Tower-Struggle. Additionally, SlowFast-R50 [16] outperforms TSN [68] in both struggle label distribution regression and classification, showcasing that 3D-ConvNets is more advanced in extracting spatial-temporal features than the 2D-ConvNets. VTN-SlowFast-ViT further surpasses SlowFast-R50 [16] in general, suggesting that its additional ViT layer helps better align predictions with the voters’ label distributions.
@@ 6 Attention Visualization
In this section, we visualize what the models focus on using attention weights and activation maps [5], ensuring they effectively detect participants’ struggles. Specifically, we examine activation maps from the backbone network and attention scores from the self-attention layer of the VTN-SlowFast-ViT model.
Figure 8 presents heatmaps from the final convolutional layer of the backbone network, generated using Grad-CAM++ [5], along with attention scores visualized as green bars. We provide examples from three activities: Pipes-Struggle, Tent-Struggle, and Tower-Struggle. The heatmaps reveal that the model predominantly focuses on hands and active objects, though occasionally not robust enough, causing attention to drift toward the background. For example, in the Pipes-Struggle visualised frames, the heatmaps highlight the hand of the person holding the pipes as well as the pipes that are lying in the background. Similarly, in the Tent-Struggle examples, the model focuses on the tent’s poles and the hands of the person attempting to put it up, as shown by the heatmaps. In the Tower-Struggle examples, the heatmaps mainly lie on the hands and the Hanoi puzzle blocks. The corresponding green bars below show that the model is giving each frame varying amounts of attention. Since the analysed video clips are limited to 10 seconds and do not capture the full action sequence, the green bars may not align perfectly with explicit actions. However, they still indicate the model’s preference for specific frames within each sequence.
7 Limitations
We acknowledge several limitations and areas for improvement as follows: (1) Our dataset primarily consists of two indoor environments for pipe assembly and the Tower of Hanoi game, along with one outdoor setting for tent assembly. This may limit its generalizability to more diverse backgrounds and is also constrained by the amount of available data. (2) Struggle in our dataset is annotated based on 10-second trimmed video segments, which may limit the ability to achieve finer-grained temporal localization and capture struggle over longer temporal dependencies.
@@ 8 Discussion and Conclusions
In this paper, we introduced the struggle determination task as an important component in video understanding. We collected a new dataset for struggle determination, comprising three daily activities–Pipes-Struggle, Tent-Struggle, and Tower-Struggle–with a total of 5.09 hours of video recordings. Struggle annotations were provided at the video level by both human experts and crowd voters.
Our experiments evaluated various deep learning models as baselines for struggle determination. Among the different modelling approaches, struggle classification proved to be the most effective. The struggle classification results reveal opportunities for improving four-way struggle classification, which particularly focuses on finer-grained struggle assessment. The similar accuracy between four-way to binary conversion and binary classification suggests that the main challenge lies in distinguishing different degrees of struggle—an essential factor for developing reliable assistive systems.
We further conducted struggle generalization experiments across the activities showcasing the challenges, and we also compared different models to identify the most efficient backbone model architectures for struggle determination. Our findings in the ablation study highlight the importance of temporal motion information and the correct frame sequence for struggle determination. Additionally, we report baseline results for alternative training strategies, treating struggle determination as regression and label distribution learning tasks.
In conclusion, our research work in this paper lays the foundation for further research in struggle determination from videos. We hope this research provides valuable insights for additional work on fine-grained egocentric video understanding and supports new methods and directions to benchmark assistive systems.
@@
Dataset Access Our struggle determination dataset can be found via this URL: https://github.com/FELIXFENG2019/Struggle-Determination.
@@
Acknowledgments We are grateful to Prof. Dima Damen’s contributions to early discussions and involvement in the initial investigation of this work. Shijia Feng is supported by a scholarship from the China Scholarship Council (No.202109210007). Initial research, data capture and annotation supported by EPSRC grant GLANCE (EP/N013964/1).
@@ Appendix A Related Datasets
In comparison, our struggle determination datasets encompass three tasks that demonstrate different types of skills in assembling pipes, pitching tents, and playing the Tower of Hanoi game. These datasets cover indoor and outdoor scenes and include both rigid and non-rigid objects. Additionally, we specifically use a struggle scoring scale. Thus, the annotations in our dataset are not limited to binary labels but are further fine-grained to indicate the degree of struggle. Aside from the focus on struggle, this distinguishes them from the class labels in AR, action quality scores in AQA, ranking labels in skill determination, and even the binary contrastive labels in VAD. Hence, it is essential to create a dataset with a focus on struggle annotations to train deep models for determining struggle and quantifying its degree.
@@ B.1 Instructions for Participants
During our data collection process, instructions were given to our in-person participants both verbally and written on paper. For the plumbing pipes task, there are two instruction diagrams illustrating how to assemble the pipes correctly. One version is easy to follow, shown in Figure 9, and another more difficult version, shown in Figure 10 so that the participants using a hard version of instructions are more likely to show signs of struggle. For the tent-pitching task, we used the paper instructions from the manufacturer, as shown in Figure 11. For the Tower of Hanoi game task, the researchers only provided the instructions verbally introducing the basic rules of the game, such as only one disk can be moved at a time, and a larger disk cannot be placed on top of a smaller one.